A rectangular parking lot is 50 ft longer than it is wide. Determne the dimensions of the parking lot if it measures 250 ft. diagonally.
Answer by dolly(163)

Let the width of the parking lot be 'x' ft
So length = (x+50) ft
Given the diagonal measures = 250 ft
The width, length and the diagonal together form a right triangle.
Applying Pythagorean theorem,
x^2 + (x+50)^2 = (250)^2
==> x^2 + x^2 + 100x + 2500 = 62500
==> 2x^2 + 100x + 2500 - 62500 = 0
==> 2x^2 + 100x - 60000 = 0
==> x^2 + 50x - 30000 = 0 [dividing by 2]
==> x^2 + 200x - 150x - 30000 = 0
==> x(x+200) - 150(x + 200) = 0
==> (x+200) (x-150) = 0
==> x+200 = 0 or x-150 = 0
As x denotes width, x cannot be negative.
Thus we take x = 150 only
So the length of the parking lot = x+50 = 200 ft
Width of the parking lot = x = 150 ft.